OpenStudy (anonymous):

what is the formula for findig the inverse secant of 2/sqrt3

OpenStudy (anonymous):

let y= sec^-1 (2/ sqrt(3))

OpenStudy (anonymous):

therefore sec(y) = 2/sqrt(3)

OpenStudy (anonymous):

but sec = 1/cos

OpenStudy (anonymous):

so cosy = sqrt(3) / 2

OpenStudy (anonymous):

so y= pi/ 6 , which was what we were asked to find, we let y be what we wanted